# Defending against Indirect Prompt Injection by Instruction Detection

### v1

@@ Results
Ours 99.60% 98.35% 99.45% 96.90%
Table 3: Comparison of detection accuracy between solely utilizing hidden states, solely utilizing gradients, and our proposed method combining hidden states and gradients. The highest detection accuracy is indicated in bold. Here, ID denotes the in-domain setting, whereas OOD denotes the out-of-domain setting.
Detection Accuracy across Different Layers
@@
We further examine the detection accuracy of solely utilizing hidden states or gradients across different layers on all dataset combinations. As presented in Figure 2, the detection accuracy across different layers demonstrates a clear trend: performance initially improves with increasing layer depth, reaches a peak at the middle layers, but then fluctuates significantly and generally declines. This trend highlights that intermediate layers capture more informative features relevant to instruction detection, whereas deeper layers may introduce noise or less task-specific representations, which is consistent with our observations on the validation set. These findings also align with the observations of recent study [41], demonstrating that intermediate layers in LLMs often yield richer representations for downstream tasks compared to the final layers.
Large Language Models
@@
The effectiveness of our proposed method is evaluated across various LLMs, including different architectures (Llama [38], Qwen [42], Mistral [43]) and model sizes (1B, 3B, 7B, 8B, 14B parameters). As shown in Table 4, features extracted from all tested LLMs are effective in detecting hidden instructions. Notably, we select the hidden states and gradients from the best-performing layer, which are all located in the intermediate layers. Among the evaluated models, Qwen-2.5-7B and Llama-3.1-8B exhibit superior results, while Mistral-7B shows slightly less optimal performance. Furthermore, the findings indicate that larger models generally produce features that are more effective for instruction detection, aligning with our hypothesis that stronger model capabilities lead to features that better facilitate the identification of hidden instructions.
@@
We also include a comparison between Llama-3.1-8B-Base and Llama-3.1-8B-Instruct. The results show a significant performance gap, with Llama-3.1-8B-Base demonstrating notably worse results. This difference is likely due to the fact that the success of IPI attacks relies on the presence of hidden instructions embedded within external content, which alter the behavioral state of LLMs. Since Llama-3.1-8B-Base has not undergone instruction fine-tuning, it does not exhibit the same responsiveness to such hidden instructions in the way that the instruct model does. As a result, the ability of Llama-3.1-8B-Base to detect such attacks is considerably diminished.
Training Data Size
@@
Figure 3: Comparison of detection accuracy between LLM fine-tuning and our proposed method on different training data size. (a) Detection accuracy comparison in the in-domain setting (Wikipedia+LaMini-Instruction). (b) Detection accuracy comparison in the out-of-domain setting (News Article+BIPIA).
@@
We conduct experiments using training data of varying sizes to assess how the quantity of training data affects the performance of our method. As presented in Figure 3, even with a small training set of only 50 samples (25 positive and 25 negative), our method achieves relatively high performance, exceeding 95% accuracy in both in-domain and out-of-domain scenarios. These results indicate that our approach requires only minimal training data to achieve strong results. These findings highlight the remarkable data efficiency of our method, which performs well even with very limited data.
Composition of Training Data
@@
We conduct experiments using different combinations of training datasets to assess the robustness and adaptability of our method to various training dataset compositions. As presented in Table 5, our approach consistently yields high accuracy across all test datasets, regardless of the specific combination of training data used. This indicates the generalizability and adaptability of our approach, as it does not rely on any particular training dataset source. Additionally, we observe that accuracy is consistently lower when tested on the News Articles with the BIPIA combination, indicating that this scenario poses the greatest challenge for instruction detection. Nonetheless, our method still achieves satisfactory accuracy in this challenging scenario, further validating its effectiveness and robustness in instruction detection.
Paired Responses for Gradients
@@
To investigate the effect of various paired responses on the extraction of gradient features, we conduct experiments using four candidate responses: "I’m sorry" "Hello" "Yes" and "Sure." These candidates are selected based on an analysis of common responses to instructions in WildChat [44] dataset, ranked by frequency. Results in Table 6 show that all four paired responses achieve high accuracy (>95%) in distinguishing between normal external data and those containing hidden instructions. Among them, "Sure" delivers the best performance across all test datasets, further validating our choice of "Sure" as the paired response in our method. These results emphasize the robustness of our approach to differentiate response pairings while confirming that "Sure" is a particularly effective option for this task.
Discussion
As LLMs are increasingly adopted in real-world applications, the inclusion of external data provides significant advantages but also introduces new security challenges. A particularly concerning threat is IPI, where adversaries embed covert instructions within external data to manipulate LLMs into performing unintended or potentially harmful actions. Mitigating this vulnerability is not just a theoretical issue, but a practical necessity for ensuring the safety, reliability, and trustworthiness of LLM-integrated systems. As LLMs are increasingly deployed in sensitive environments—from healthcare to finance and law—the ability to proactively detect and mitigate such threats is crucial for guaranteeing the safe and responsible deployment of AI systems.
Our work proposes a detection-based approach that leverages the internal behavioral states of LLMs as signals to identify IPI attacks. A key strength of our method lies in its task-agnostic design. Unlike existing approaches that rely on runtime interventions, our method enables proactive screening of external sources, allowing the removal of external data containing covert instructions in advance. This eliminates the need for additional steps during task execution, minimizing time overhead and ensuring a smoother, more efficient user experience. Another unique aspect of our approach is its use of internal behavioral signals as indicators of covert instruction presence. This design aligns closely with the fundamental mechanism of IPI attacks: since IPI attacks fundamentally rely on the external instructions to alter the behavioral state of LLMs, and our approach leverages the distinct behavioral states of LLMs to differentiate between data and instructions, it becomes inherently difficult for attackers to evade detection without compromising the effectiveness of the attack. Moreover, by exploiting the LLM’s natural sensitivity to embedded instructions, our method achieves high detection performance with only a small number of labeled samples, making it both robust and data-efficient. We believe our work can spark further research into IPI attacks in LLM-integrated systems. Moreover, we hope it inspires broader efforts to secure LLM-integrated systems, especially in sensitive domains such as healthcare, law and finance, where the risks of model manipulation can be severe.
Nevertheless, our method has several limitations. First, it requires both forward and backward passes through the LLM, introducing additional computational overhead compared to lightweight defenses. While suitable for offline filtering or batch processing, this may limit deployment in resource-constrained settings. Second, although our experiments cover multiple representative scenarios and datasets, we cannot guarantee coverage of all possible attack strategies or domain-specific variations. Third, the current design adopts a conservative binary decision—discarding any external data flagged as containing hidden instructions—result in the unintended removal of useful, non-malicious information. In future work, when hidden instructions are identified, we will attempt to refine this approach by isolating and eliminating the hidden instructions embedded within the external data, rather than discarding the entire external data. This enhancement could enable the LLMs to leverage the remaining valid information while maintaining robust defenses against hidden instructions.
Ethical and Societal Impact
Our proposed method defends against IPI attacks, which is essential for ensuring the secure and reliable operation of LLMs in third-party system integrations. By mitigating the risks posed by IPI attacks, our approach fosters ethical and socially responsible use of AI technologies, enhancing trust in their application within critical sectors such as healthcare, legal and finance domains. There may be concerns about whether our method could provide attackers with insights to bypass detection. Since our approach leverages the distinct behavioral states of LLMs to differentiate between data and instructions, while IPI attacks fundamentally rely on the external instructions to alter the behavioral state of LLMs, it would be exceedingly difficult for attackers to circumvent our detection. In summary, our method strengthens the security and trustworthiness of AI systems by effectively defending IPI attacks, aligning with ethical principles and supporting the development of reliable, safe, and socially responsible AI technologies for real-world applications.
Method
Related Work
Indirect Prompt Injection Defense
Defending against IPI attacks is a critical research area in ensuring the secure and reliable use of LLMs [17, 18, 19]. Existing defenses are generally classified into prevention-based defences·and detection-based defences [27, 28].
Prevention-based defenses primarily focus on ensuring LLMs to follow user instructions while ignoring external ones. These approaches are further divided into black-box defenses and white-box defenses. Black-box defenses [27, 45, 46, 47] typically aim to isolate user instructions from external data, using carefully designed prompts to ensure LLMs disregard any hidden instructions within the external data. These methods work without access to the internal parameters of the model, focusing on input preprocessing and separation mechanisms. In contrast, white-box defenses [27, 48] utilize the internal parameters of the model and involve fine-tuning LLMs with samples of IPI attacks. By training on a diverse set of IPI scenarios, these methods enhance the robustness of LLMs to ignore external instructions while maintaining performance on the intended task.
Detection-based defenses, though relatively underexplored, aim to identify IPI attacks and can be generally divided into three main strategies. LLM (Zero-shot) [28] directly uses LLMs to identify hidden instructions in external data. Response Check [28] evaluates whether the model’s outputs remain consistent with the intended task. TaskTracker [36] detects IPI attacks by contrasting the LLM’s activations before and after feeding the external data, which indicates whether the user’s instruction is distorted by the instruction hidden in the external data. Our method also falls under detection-based defenses, bridging the gap with a more robust mechanism for instruction detection.
Behavioral States of Large Language Models
Recent studies [29, 49] have delved extensively into the internal mechanisms of LLMs, identifying hidden states and gradients as highly informative features for understanding and controlling their behavior. These behavioral states are increasingly recognized for their potential to enhance the transparency, safety, and robustness of LLMs.
Hidden states, especially those from intermediate layers, have been shown to encode rich and insightful representations of given inputs. RepE [29] utilizes representations from the last token’s hidden states to monitor and manipulate high-level cognitive phenomena in LLMs. Furthermore, a recent study [41] has explored the effectiveness of intermediate features across different LLM architectures, revealing that intermediate features often yield richer information than final-layer for downstream use.
Gradients provide another critical lens for analyzing the LLM’s behavior. Gradsafe [49] leverages the observation that adversarial prompts generate distinct gradient patterns compared to safe prompts, enabling effective jailbreak prompts detection without additional training by analyzing gradients related to safety-critical parameters. Additionally, much literature [30, 31, 32] has explored the functions of self-attention layers and feed-forward layers, providing key insights into where to focus when analyzing gradients in our work. Research has shown that self-attention layers capture behavioral characteristics, such as linguistic dependencies and token relationships, while feed-forward layers encode knowledge-based features, enabling the model to leverage the knowledge learned during training. Given that the instruction recognition task primarily relies on the behavioral characteristics of LLMs, we focus on the gradients of self-attention layers.
Defending IPI attacks via Instruction-Following State Detection
Overview
In our proposed method, we aim to detect IPI attacks through the behavioral states of LLMs, hypothesizing that changes in the behavioral states of LLMs induced by embedded instructions in external content can be effectively utilized to detect such attacks. To achieve this, we fuse the hidden states and gradients from the most effective layers, integrating complementary information captured by both features. These fused features are then fed into an MLP classifier, enabling accurate and robust detection of IPI attacks. The overall framework is illustrated in Figure 1b, with detailed processes discussed in the following sections.
Hidden States Extraction
To leverage hidden states as features, we first take external data as the input of the LLM and extract the hidden states corresponding to the last token at each layer. These hidden states are then fed into an MLP classifier to assess their ability to distinguish between normal external data and those containing hidden instructions. In our approach, we use the Llama-3.1-8B-Instruct model, which consists of 32 layers. Through experimentation on the validation set, we identify that the hidden states from the 14th layer provides the best performance in instruction detection. Therefore, we select the last token’s hidden state from the 14th layer, a vector with a dimension of 4096, as the first input of the feature fusion module.
Gradients Extraction
To leverage gradients as features, we first take external data as the input of the LLM, paired with a typical response to instructions, such as "Sure," and compute the gradients for the model parameters at each layer during back propagation. Based on prior research indicating that self-attention layers capture the model’s behavioral characteristics, while the feed-forward layers are more effective at encoding knowledge-based features, we concentrate on the gradients of self-attention layers. Experimental results on the validation set demonstrate that the gradients from the 14th layer, consistent with the layer identified for hidden states, yield the best performance in distinguishing between normal external data and those containing hidden instructions.
Feature Fusion
Data Availability
The Wikipedia dataset is available at https://huggingface.co/datasets/wikimedia/wikipedia. The News Article dataset is available at https://dataverse.harvard.edu/dataset.xhtml?persistentId=doi:10.7910/DVN/GMFCTR. The LaMini-instruction dataset is available at https://huggingface.co/datasets/MBZUAI/LaMini-instruction. The BIPIA benchmark is available at https://github.com/microsoft/BIPIA.
Code Availability
The code is publicly available at https://github.com/MYVAE/Instruction-detection.
Acknowledgments
We sincerely appreciate the editor and anonymous reviewers for their insightful comments and suggestions that may significantly enhance the quality of this work.
Author contributions
T.W. conceived the idea of this work, implemented the proposed method, conducted experiments, analyzed the results, and contributed to the writing of this manuscript. C.W. was involved in the technical discussions, and contributed to the writing of this manuscript. X.Y. was involved in the technical discussions, and analyzed the results. H.T. was involved in the technical discussions, and analyzed the results. Y.X. contributed to the writing of this manuscript, and coordinated the research project. L.L. contributed to the writing of this manuscript. Z.D. was involved in the technical discussions, and contributed to the writing of this manuscript. F.W. conceived the idea of this work, analyzed the results, contributed to the writing of this manuscript, and coordinated the research project.
Additional Information
Comparison with Related Method

### v2

@@ 4.4.2 Attack Success Rate Comparison
InstructDetector 99.60% 98.35% 99.45% 96.90%
Table 3: Comparison of detection accuracy between solely utilizing hidden states, solely utilizing gradients, and InstructDetector, combining hidden states and gradients. The highest detection accuracy is indicated in bold. Here, ID denotes the in-domain setting, whereas OOD denotes the out-of-domain setting.
@@
To assess the effectiveness of InstructDetector in lowering ASR, we compare it with several prevention-based defenses, including in-context learning, multi-turn dialogue, and adversarial training. As shown in Table 2, IPI attacks remain effective against both open-access and proprietary LLMs, and in some cases even achieve higher success rates on stronger models such as GPT-4o. Nevertheless, InstructDetector consistently yields the lowest ASR across both open-access and proprietary models.
Among the baselines, in-context learning and multi-turn dialogue, which are both black-box approaches, exhibit limited effectiveness in reducing ASR on both open-access and proprietary models, with ASR remaining significantly higher than that of InstructDetector. This indicates that simple structural modifications or prompting strategies fail to provide robust protection against IPI attacks.
Adversarial training, a white-box method, demonstrates greater effectiveness in lowering ASR compared to black-box approaches. However, it still underperforms compared to our method and has limitations, especially for proprietary models, since it involves changes to the embedding layer and necessitates model fine-tuning. Our approach stands out for its ability to achieve superior ASR reduction while maintaining compatibility with both open-access and proprietary models, demonstrating its practicality and robustness against IPI attacks.
4.5 Ablation Study
For additional ablation studies, including experiments on different training data compositions, the impact of paired response, and the influence of instruction quantity and position, please refer to Appendix C.
4.5.1 Solely Utilizing Hidden States/Gradients
4.5.2 Detection Accuracy across Layers
We further examine the detection accuracy of solely utilizing hidden states or gradients across different layers on all dataset combinations. As presented in Figure 3, the detection accuracy across different layers demonstrates a clear trend: performance initially improves with increasing layer depth, reaches a peak at the middle layers, but then fluctuates significantly and generally declines. This trend highlights that intermediate layers capture more informative features relevant to instruction detection, whereas deeper layers may introduce noise or less task-specific representations, which is consistent with our observations on the validation set. These findings also align with the observations of a recent study (Skean et al., 2024), demonstrating that intermediate layers in LLMs often yield richer representations for downstream tasks compared to the final layers.
4.5.3 Large Language Models
@@
The effectiveness of InstructDetector is evaluated across various LLMs, including different architectures (Llama (Dubey et al., 2024), Qwen (Bai et al., 2023), Mistral (Jiang et al., 2023)) and model sizes (1B, 3B, 7B, 8B, 14B parameters). As shown in Table 4, features extracted from all tested LLMs are effective in detecting hidden instructions. Notably, we select the hidden states and gradients from the best-performing layer, which are all located in the intermediate layers. Among the evaluated models, Qwen-2.5-7B and Llama-3.1-8B exhibit superior results, while Mistral-7B shows slightly less optimal performance. Furthermore, the findings indicate that larger models generally produce features that are more effective for instruction detection, aligning with our hypothesis that stronger model capabilities lead to features that better facilitate the identification of hidden instructions.
@@
We also include a comparison between Llama-3.1-8B-Base and Llama-3.1-8B-Instruct. The results show a significant performance gap, with Llama-3.1-8B-Base demonstrating notably worse results. This difference is likely because IPI attacks rely on the presence of hidden instructions embedded within external content, which alter the behavioral states of LLMs. Since Llama-3.1-8B-Base has not undergone instruction fine-tuning, it does not exhibit the same responsiveness to such hidden instructions in the way that the instruct model does. As a result, the ability of Llama-3.1-8B-Base to detect such attacks is considerably diminished.
4.5.4 Training Data Size
@@
Figure 4: Comparison of detection accuracy between LLM fine-tuning and InstructDetector on different training data size. (a) Detection accuracy comparison in the in-domain setting (Wikipedia+LaMini-Instruction). (b) Detection accuracy comparison in the out-of-domain setting (News Article+BIPIA).
English 99.60% 98.35% 99.45% 96.90%
Chinese 97.40% 95.15% 96.20% 94.25%
Thai 95.75% 94.15% 93.60% 92.05%
Table 5: Detection accuracy comparison on multilingual test sets. The highest detection accuracy is indicated in bold. Here, ID denotes the in-domain setting, whereas OOD denotes the out-of-domain setting.
@@
We conduct experiments using training data of varying sizes to assess how the quantity of training data affects the performance of InstructDetector. As presented in Figure 4, even with a small training set of only 50 samples (25 positive and 25 negative), InstructDetector achieves relatively high performance, exceeding 95% accuracy in both in-domain and out-of-domain scenarios. These results indicate that InstructDetector requires only minimal training data to achieve strong results. These findings highlight the remarkable data efficiency of InstructDetector, which performs well even with very limited data.
4.5.5 Multilingual Evaluation
We construct two additional test sets to evaluate how well InstructDetector, trained solely on English data, can transfer to other languages. Specifically, following established approaches in multilingual NLP (Kassner et al., 2021), we translated our original test set using Google Translate into both a high-resource language (Chinese) and a relatively low-resource language (Thai). To ensure semantic similarity and consistency, we perform back-translation and compare the outputs with the original English versions.
As shown in Table 5, InstructDetector maintains high accuracy across these multilingual test sets, despite being trained exclusively on English data. This demonstrates that the discriminative features leveraged by InstructDetector are not language-specific, and can effectively generalize beyond English.
5 Conclusion
In this work, we present InstructDetector, a detection-based approach that leverages the internal behavioral states of LLMs as signals to identify IPI attacks. A key finding of our study is that the hidden states and gradients from the intermediate layers of LLMs provide highly discriminative features for instruction detection. By leveraging these internal behavioral states, InstructDetector provides a robust mechanism for identifying hidden instructions within external data.
We demonstrate that InstructDetector achieves superior detection accuracy in both in-domain and out-of-domain settings, while significantly reducing the attack success rate compared to existing defense methods. These findings underline the effectiveness and adaptability of InstructDetector, offering a robust solution for enhancing the security of LLM-based systems.
6 Limitation
InstructDetector has several limitations. First, it requires both forward and backward passes through the LLM, introducing additional computational overhead compared to lightweight defenses. While suitable for offline filtering or batch processing, this may limit deployment in resource-constrained settings and raise environmental concerns due to the increased energy consumption associated with higher computational demands. Second, although our experiments cover multiple representative scenarios and datasets, we cannot guarantee coverage of all possible attack strategies or domain-specific variations. Third, the current design adopts a conservative binary decision—discarding any external data flagged as containing hidden instructions—resulting in the unintended removal of useful, non-malicious information. However, the high accuracy of our method forms a solid foundation for precisely identifying and removing external instructions. By recursively segmenting documents and applying our method to each segment individually, we can remove only indirect prompts while preserving useful content.
7 Ethical Impact
Our proposed method, InstructDetector, defends against IPI attacks, which is essential for ensuring the secure and reliable operation of LLMs in third-party system integrations. By mitigating the risks posed by IPI attacks, InstructDetector fosters ethical and socially responsible use of AI technologies, enhancing trust in their application within critical sectors such as healthcare, legal, and finance domains. There may be concerns about whether InstructDetector could provide attackers with insights to bypass detection. Since InstructDetector leverages the distinct behavioral states of LLMs to differentiate between data and instructions, while IPI attacks fundamentally rely on the external instructions to alter the behavioral states of LLMs, it would be exceedingly difficult for attackers to circumvent our detection. In summary, InstructDetector strengthens the security and trustworthiness of AI systems by effectively defending IPI attacks, aligning with ethical principles, and supporting the development of reliable, safe, and socially responsible AI technologies for real-world applications.
8 Acknowledgments
We sincerely appreciate the anonymous reviewers for their insightful comments and suggestions that may significantly enhance the quality of this work. The work was partially done at the Beijing Key Laboratory of Research on Large Models and Intelligent Governance.
Appendix A Additional Experimental Settings
Prompt for LLM (Zero-shot). ### Guideline ### You are a helpful assistant, and your task is to determine if there is any external instruction in the text. Respond with 1 if there is an instruction, or 0 if there is not. Only reply with 1 or 0, without additional explanation. ### Your Task ### Text: external data Answer:
Prompt for LLM (Few-shot). ### Guideline ### You are a helpful assistant, and your task is to determine if there is any external instruction in the text. Respond with 1 if there is an instruction, or 0 if there is not. Only reply with 1 or 0, without additional explanation. ### Examples ### Here are a few concrete examples to help you understand how to make a decision. Example 1: Text: external data Answer: label ### Your Task ### Text: external data Answer:
Prompt for Response Check. Please follow the instruction based on the context below. Instruction: instruction Context: external data
@@ Appendix A Additional Experimental Settings
A.1 Dataset Details
All datasets used in this work are in English, and the use of these datasets complies with their intended purposes as specified by their creators. We have carefully reviewed the data to ensure that they do not contain personally identifiable information or offensive content.
Wikipedia
The dataset is constructed using Wikipedia dump files, under the CC-BY-SA license. Each data instance comprises the content of an entire Wikipedia article. In addition, we remove the overly long articles to ensure that they are not truncated during processing.
News Articles
The dataset contains 3,824 news articles, each featuring metadata including the title, subtitle, content, and publication date, sourced from multiple media outlets, under the CC0 license. Similarly, we remove the overly long articles to ensure that they are not truncated during processing.
LaMini-instruction
The dataset consists of 2.58 million pairs of instructions and corresponding responses, generated using GPT-3.5-Turbo, drawing from a wide range of existing resources of prompts, including Self-Instruct, P3, FLAN, and Alpaca, under the CC-BY-NC license.
BIPIA
BIPIA is the first benchmark aimed at evaluating the risk of IPI attacks on LLMs, under the MIT license, and we use its instruction dataset for our experiments. The dataset consists of 15 attack types, categorized into task-irrelevant, task-relevant, and targeted attacks, with 5 instructions per attack type, resulting in a total of 75 instructions across both the training and test sets. These instructions were semi-automatically generated with the assistance of ChatGPT and manually reviewed for rationality.
A.2 Baseline Details
A.2.1 Detection-based Defenses
LLM (Zero-shot) (Liu et al., 2024)
Response Check (Liu et al., 2024)
TaskTracker (Abdelnabi et al., 2024)
LLM (Few-shot)
To enhance the performance of Naive LLM-based Detection, we attempt to leverage in-context learning (Dong et al., 2022) to strengthen the model’s capability to detect hidden instructions, where task demonstrations are integrated into the textual prompt.
LLM (Fine-tuning)
A.2.2 Prevention-based Defenses
In-context Learning (Yi et al., 2025)
Employ in-context learning to enable the model to distinguish between external data and user instructions, by providing samples where the model responds to input containing external data without being misled by the instruction embedded within external data.
Multi-turn Dialogue (Yi et al., 2025)
Adversarial Training (Yi et al., 2025)
Incorporate adversarial learning during the LLM’s self-supervised fine-tuning phase, training the model to disregard instructions embedded within external content. The approach further adapts the model’s embedding layer to explicitly demarcate external content boundaries, enabling clearer distinction between external content and user instructions.
A.3 Additional Experimental Details
As shown earlier, we employ specific prompt templates for LLM (Zero-shot), Response Check, and LLM (Few-shot).
For the LLM (Fine-tuning), we use the AdamW optimizer to train the model for one epoch, with a learning rate set at 0.00001 and a maximum input length of 2048 tokens. Both the learning rate and the number of epochs are chosen based on performance on the validation set. For TaskTracker, we use the publicly available model released by the work. As for the prevention-based approaches, we utilize the publicly available implementations provided by the respective works.
We conduct our experiments using an 8B-parameter LLM. All training and inference are performed on a single NVIDIA A100 80GB GPU, with the training phase requiring approximately 1 GPU hour. In practice, InstructDetector can process 1283 tokens per second on a single A100 GPU. It is also compatible with various acceleration and quantization techniques, as long as gradients and hidden states are accessible, which can further improve efficiency. Importantly, unlike existing task-specific approaches that require runtime interventions, InstructDetector allows for pre-screening of external sources and caching of results. This design significantly reduces the latency incurred during actual system usage and makes the overhead of instruction detection negligible at inference time. For all reported results, we present outcomes from a single run. During preprocessing, to construct negative samples by inserting instructions into benign text, we utilize the PunktSentenceTokenizer from the NLTK library for sentence segmentation.